OpenStudy (anonymous):

Simplify 5a(a – 7) 5a2 – 35a 5a2 + 35 5a2 – 7 5a2 – 7a

Parth (parthkohli):

Hey! Do you know how to use the distributive property?

Parth (parthkohli):

\( \color{Black}{\Rightarrow a(b \pm c) = ab \pm ac \Longrightarrow 5a(a - 7) = (5a \times a) - (5a \times 7) }\)

OpenStudy (anonymous):

i dont kno the answer

OpenStudy (unklerhaukus):

\[5a(a – 7)=(5a\times a) + (5a\times-7)\]

OpenStudy (anonymous):

Ansere is A..

OpenStudy (anonymous):

5a * a + 5*a*(-7) = 5a2 - 35a... * means Multiply...